# A tuning-free and scalable method for joint graphical model estimation with sharper bounds

### v2

@@ 3.2 Support recovery and asymptotic distribution
Assumption 4 (Minimum signal conditions)
There exists a sufficiently large constant $C_{1}>0$ such that
$\min_{j\in[p],k\in[K]}\min_{i\in S_{j}^{(k)}}\left|\Theta^{(k)}_{i,j}\right|^{2}\geq$ $\frac{C_{1}}{n_{k}}\left(\frac{\log p}{s_{0}}+\log(sK)\right),$
$\min_{j\in[p]}\min_{i\in S_{j}}\sum_{k\in[K]}\frac{n_{k}}{N}\left(\Theta^{(k)}_{i,j}\right)^{2}\geq$ $C_{1}\left(\frac{\log p+s_{0}\log(sK)}{N}\right).$
Theorem 2 establishes the sufficiency of the above signal conditions for selection consistency.
@@
Assume Assumptions 1-4 hold and $n\gtrsim K+s\log p+ss_{0}\log(sK)$. Assume that there exists a constant $\gamma\in(0,1)$ such that $\frac{\log p}{s_{0}}+\log(sK)\geq\gamma\log p$. Then we simultaneously have the following properties with probability at least $1-C_{1}e^{-C_{2,\gamma}\left(\frac{\log p}{s_{0}}+\log(sK)\right)}$:
@@
    Selection consistency:
    $\left(\hat{S}_{j}^{(1)},\cdots,\hat{S}_{j}^{(K)}\right)=\left(S_{j}^{(1)},\cdots,S_{j}^{(K)}\right)\text{ for every }j\in[p].$
@@
    Sharper estimation error bound:
    $\max_{j\in[p]}\sum_{k\in[K]}\frac{n_{k}}{N}\left\|\hat{\Theta}^{(k)}_{\cdot,j}-\Theta^{(k)}_{\cdot,j}\right\|_{2}^{2}\lesssim\frac{K+ss_{0}+\log p}{N}.$
Building on the framework of Wang et al. (2010), we can also establish the necessity of the signal condition required in Theorem 2. This implies that no algorithm can guarantee such exact support recovery if Assumption 4 is violated (in terms of rate). Thus our method achieves rate optimality in terms of selection consistency. Additionally, Theorem 2 shows that our method is inherently signal-adaptive, achieving enhanced estimation accuracy under proper signal conditions. This is not found in other joint estimation methods.
Define $\hat{S}_{j+}^{(k)}:=\hat{S}_{j}^{(k)}\cup\{j\}=\text{supp}\left(\hat{\Theta}_{\cdot,j}^{(k)}\right)$. The following theorem establishes the asymptotic property of our method.
@@
Assume that Assumptions 1-4 hold, and that $\sqrt{n}\succ K+s\log p+ss_{0}\log(sK)$. Additionally, assume there exists a constant $\gamma\in(0,1)$ such that $\frac{\log p}{s_{0}}+\log(sK)\geq\gamma\log p$. Then, as $n\to\infty$, we have
$\max_{k\in[K],j\in[p]}~{}\max_{i\in\hat{S}_{j+}^{(k)}}~{}\sup_{w\in\mathbb{R}}\left|\mathbf{P}\left(\frac{\sqrt{n_{k}}\left(\hat{\Theta}^{(k)}_{i,j}-\Theta^{(k)}_{i,j}\right)}{\sigma_{i,\hat{S}_{j}}^{(k)}}\leq w\right)-\Phi(w)\right|\to 0,$
@@
Theorem 3 demonstrates that our MIGHT method possesses asymptotic normality, thereby enabling statistical inference. This property distinguishes our method from existing joint estimation methods. In application, we estimate $\left(\sigma^{(k)}_{i,\hat{S}_{j}}\right)^{2}$ by using
$\left(\hat{\sigma}^{(k)}_{i,\hat{S}_{j}}\right)^{2}=\frac{1}{n_{k}}\sum_{\ell=1}^{n_{k}}\left\{\left(\hat{\Sigma}^{(k)}_{\hat{S}_{j+}^{(k)},\hat{S}_{j+}^{(k)}}\right)^{-1}_{i,\cdot}X_{\ell,\hat{S}_{j+}^{(k)}}^{(k)}\left(X_{\ell,\hat{S}_{j+}^{(k)}}^{(k)}\right)^{\top}\left(\hat{\Sigma}^{(k)}_{\hat{S}_{j+}^{(k)},\hat{S}_{j+}^{(k)}}\right)^{-1}_{\cdot,j}\right\}^{2}-\left(\hat{\Theta}_{i,j}^{(k)}\right)^{2},$
@@ 4 Numerical experiments
This section evaluates the performance of our proposed method, MIGHT, against five existing methods: three joint estimation methods: the joint estimator method by Guo et al. (2011) (JEM), the group graphical lasso by Danaher et al. (2014) (GGL), and the fast and scalable joint estimator by Wang et al. (2017) (FJEM); and two separate estimation methods: the separate graphical lasso by Friedman et al. (2008) (Sep Glasso) and the separate nodewise $L_{0}$-regression by Shu et al. (2024) (Sep Node). We use the Bayesian Information Criterion (BIC) to select the optimal hyperparameters for these five methods.
We set $p=100,K=10$ and $n_{1}=\cdots=n_{K}=100$. Data generation begins with an Erdős–Rćenyi base graph $G$, where each pair of nodes connects independently with a probability of $0.1$. We generate $K$ graphs $G_{1},\cdots,G_{K}$ by independently pruning a fraction $\rho$ of edges from the base graph $G$, with $\rho\in(0,1)$ controlling neighbor-similarity and edge-sparsity. The value of each edge in $G_{k}$ is uniformly sampled from $[-1,-0.5]\cup[0.5,1]$, forming the adjacency matrix $\Omega^{(k)}$. Then, the positive-definite precision matrix $\Theta^{(k)}$ is constructed as $\Theta^{(k)}=\Omega^{(k)}+\left(r+|\Lambda_{\min}(\Omega^{(k)})|\right)\cdot\mathbf{I}_{p}$, where parameter $r>0$ modulates the signal strength by inflating diagonal entries, thereby reducing the signal-to-noise ratio (see (3)) in the nodewise regression. Finally, by applying the corresponding covariance matrices $\left\{\Sigma^{(k)}\right\}_{k\in[K]}=\left\{(\Theta^{(k)})^{-1}\right\}_{k\in[K]}$, we get $K$ datasets as $X^{(1)},\cdots,X^{(K)}$.
@@ 4.1 Simulation 1: similarity and sparsity
We fix $r=0.1$ and vary $\rho\in\{0.2,0.5,0.8\}$, with each setting 100 times simulation. Table 1 presents the simulation results; all MCC values have been multiplied by 100, and standard errors are shown in parentheses. Our MIGHT method demonstrates superior performance across all evaluation metrics compared to competing methods. As $\rho$ increases, the recurrence frequency of shared neighbors across the $K$ graphical models decreases, leading to diminished structural similarity and a decline in the MCC-Ngbr metric.
@@
$\rho$ Methods Frobenius Norm Max L2 Norm MCC-Edge MCC-Ngbr
0.2 MIGHT 18.979 (0.113) 6.257 (0.088) 63.662 (0.305) 89.673 (0.267)
@@ 4.2 Simulation 2: signal strength
We fix $\rho=0.5$ and vary $r\in\{0.2,0.25,0.33,0.5,1\}$, with $1/r\in\{1,2,3,4,5\}$ measuring the signal strength. The results are presented in Figure 2. As the signal strength increases, the MIGHT method demonstrates lower estimation error and higher identification accuracy, which confirms that the enhanced signal strength contributes to better performance, consistent with our theoretical results.
Figure 2: Performance metrics with increasing signal strength ($1/r$), with each point $50$ times repetitions. All MCC values have been multiplied by 100. FJEM exhibits unstable estimation and we exclude its excessively large estimation errors.
@@ 4.3 Simulation 3: asymptotic distribution
We fix $p=50,\rho=0.5,r=0.1$ and compare the asymptotic performance of MIGHT, GGL, JEM, and Sep Node methods with two reference cases $\Theta_{2,3}^{(1)}$ and $\Theta_{5,6}^{(2)}$. The results in Figure 3 demonstrate that the MIGHT method achieves better asymptotic normality compared to the separate $\ell_{0}$ estimator Sep Node, while GGL and JEM, two joint estimation methods based on M-estimator, fail to exhibit this property. More reference cases are provided in the supplementary material.
Figure 3: Histogram of $\sqrt{n}\left(\hat{\Theta}_{ij}^{(k)}-\Theta_{ij}^{(k)}\right)/\hat{\sigma}_{ij}^{(k)}$ among different methods with $300$ times repetitions. The orange curve represents the density of the standard normal distribution, while the black and orange vertical lines indicate the empirical mean and the population mean (0), respectively.
@@ 5 Real data analysis
In this section, we apply our MIGHT method to a gene-expression cancer RNA-seq dataset, which is accessible via the UCI Machine Learning Repository (Fiorini, 2016). This dataset encompasses 801 patients and comprises 20,531 genes. The patients are categorized into five distinct types based on various tumor types, namely BReast CAncer (BRCA), KIdney Renal cell Carcinoma (KIRC), COlon ADenocarcinoma (COAD), LUng ADenocarcinoma (LUAD), and PRostate ADenocarcinoma (PRAD). The numbers of patients in these subtypes are 300, 78, 146, 141, and 136, respectively. Although these subtypes exhibit biological differences, they all involve tumor tissue, which necessitates the joint graphical model estimation.
@@ 5.2 The classification tasks based on QDA
$\delta_{k}(\mathbf{x})=\log\hat{\pi}_{k}+\frac{1}{2}\log\left|\hat{\Theta}^{(k)}\right|-\frac{1}{2}(\mathbf{x}-\hat{{\mu}}^{(k)})^{\top}\hat{\Theta}^{(k)}(\mathbf{x}-\hat{{\mu}}^{(k)}),\text{ for each }k=1,\cdots,5,$ (8)
where $\hat{\pi}_{k}=n_{k}/N$ is the proportion of observations belonging to the class $k$, ${\hat{\mu}}_{k}=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\bm{x}_{i}^{(k)}$ is the empirical mean for the training set of each class $k$. The classification rule is then given by $\hat{k}(\mathbf{x})=\arg\max_{k=1,\cdots,5}\delta_{k}(\mathbf{x})$.
@@ 6 Conclusion and discussion
This paper proposes a tuning-free and parallelizable method for the joint estimation of multiple graphical models. Theoretical analysis demonstrates that the proposed method achieves superior estimation accuracy and establishes selection consistency and asymptotic property. Both simulation studies and empirical analyses validate the superiority of our method.
@@
In addition, our proposed joint estimation framework may offer advantages in cutting-edge applications. For instance, in the cross-domain recommendation task, each user’s behaviors from different domains regarding the same items are usually relevant. Joint estimation can effectively identify these similar structures across domains, enabling cross-domain knowledge fusion through multi-graph neural networks (Ouyang et al., 2021). Moreover, our framework shows broader applicability in multiple directed acyclic graphs, multimodal data analysis, and so on. We leave them for future exploration.
Appendix A Node-specific similarity and sparsity
In Section 2, we define the neighbor-similarity and edge-sparsity of the joint graphical model by taking the maximum over all nodes. However, in certain cases, it may be more suitable and flexible to consider node-specific similarity and sparsity parameters $(s_{0,j},s_{j})$ for each node $j\in[p]$. Notably, our tuning-free algorithm adaptively accommodates such heterogeneity among nodes, achieving a column-wise $\ell_{2}$ error bound as
$\sum_{k=1}^{K}\frac{n_{k}}{N}\left\|\hat{\Theta}^{(k)}_{\cdot j}-\Theta^{(k)}_{\cdot j}\right\|_{2}^{2}\lesssim\frac{K+s_{j}\log p+s_{j}s_{0,j}\log(s_{j}K)}{N},\text{ for each }j\in[p],$
without any structural modifications, see the proof of Theorem 1 for more details.
Appendix B Additional simulation results
This section demonstrates the additional experiment results with $p=50$, and the other parameter settings are the same as in Section 4.
We first analyze the influence of node-similarity and edge-sparsity on the joint estimation. We fix $r=0.1$ and vary $\rho\in\{0.2,0.5,0.8\}$, with each setting 100 times replication. Table 3 presents the simulation results; all MCC values have been multiplied by 100, and standard errors are shown in parentheses. It shows that our MIGHT method demonstrates superior performance across most evaluation metrics. We find that an increasing $\rho$ leads to a diminishing structural similarity, and thus a decline in the MCC-Ngbr metric.
Table 3: The estimation performance across varying $\rho$ with $p=50$.
$\rho$ Methods Frobenius Max L2 MCC-Edge MCC-Ngbr
0.2 MIGHT 6.951 (0.067) 1.928 (0.045) 83.751 (0.280) 98.699 (0.093)
GGL 9.201 (0.076) 3.001 (0.053) 57.190 (0.230) 30.577 (0.710)
JEM 8.782 (0.096) 2.672 (0.059) 87.597 (0.399) 96.549 (0.267)
FJEM 60.033 (1.600) 653.368 (131.042) 6.829 (0.396) 0.057 (0.023)
Sep Glasso 13.275 (0.085) 5.475 (0.074) 53.101 (0.212) 33.807 (0.301)
Sep Node 10.377 (0.072) 4.148 (0.081) 45.675 (0.382) 89.364 (0.355)
0.5 MIGHT 4.866 (0.045) 1.026 (0.022) 88.946 (0.224) 96.838 (0.133)
GGL 6.800 (0.047) 1.646 (0.025) 59.533 (0.344) 48.588 (1.163)
JEM 6.292 (0.063) 1.411 (0.027) 85.810 (0.232) 96.443 (0.179)
FJEM 33.844 (0.590) 106.018 (10.380) 8.919 (0.558) 0.241 (0.055)
Sep Glasso 9.944 (0.059) 3.164 (0.037) 57.033 (0.150) 36.954 (0.198)
Sep Node 7.258 (0.054) 2.034 (0.035) 59.995 (0.356) 90.345 (0.270)
0.8 MIGHT 2.758 (0.020) 0.378 (0.008) 92.413 (0.158) 94.176 (0.174)
GGL 3.701 (0.026) 0.535 (0.009) 42.573 (0.326) 45.613 (0.368)
JEM 3.636 (0.029) 0.524 (0.010) 79.929 (0.173) 92.781 (0.217)
FJEM 10.526 (0.274) 11.719 (0.881) 15.346 (1.055) 6.305 (0.677)
Sep Glasso 6.050 (0.029) 1.290 (0.016) 58.560 (0.134) 45.310 (0.168)
Sep Node 3.667 (0.026) 0.575 (0.011) 78.656 (0.272) 83.122 (0.289)
We next analyze the influence of signal strength on the joint estimation. We fix $\rho=0.5$ and vary $r\in\{0.42,0.51,0.67,0.95,1.67\}$, with $1/r\in\{0.6,1.05,1.5,1.95,2.4\}$ measuring the signal strength. The results are presented in Figure 5. Under varying signal strengths, our MIGHT method consistently achieves the lowest estimation error and highest identification accuracy.
Figure 5: Performance metrics with increasing signal strength ($1/r$) and $p=50$, with each point $50$ times repetitions. FJEM exhibits unstable estimation with high signal strength, therefore we exclude some of the excessively large estimation errors obtained by this method.
We finally show the asymptotic performance of MIGHT, GGL, JEM, and Sep Node methods. We fix $p=50,~{}\rho=0.5,~{}r=0.1$ and demonstrate the asymptotic properties of more entries $\hat{\Theta}_{ij}^{(k)}$ with $300$ times repetitions, as shown in Figure 6-9. Notably, the orange curve represents the density of the standard normal distribution, while the black and orange vertical lines indicate the empirical mean and the population mean (0), respectively. These results demonstrate that our MIGHT method achieves the strongest asymptotic normality among these four methods.
Figure 6: Histogram of (diagonal) $\sqrt{n}\left(\hat{\Theta}_{ii}^{(1)}-\Theta_{ii}^{(1)}\right)/\hat{\sigma}_{ii}^{(1)}$ among different methods.
Figure 7: Histogram of (off-diagonal) $\sqrt{n}\left(\hat{\Theta}_{ij}^{(1)}-\Theta_{ij}^{(1)}\right)/\hat{\sigma}_{ij}^{(1)}$ among different methods.
Figure 8: Histogram of (diagonal) $\sqrt{n}\left(\hat{\Theta}_{ii}^{(2)}-\Theta_{ii}^{(2)}\right)/\hat{\sigma}_{ii}^{(2)}$ among different methods.
Figure 9: Histogram of (off-diagonal) $\sqrt{n}\left(\hat{\Theta}_{ij}^{(2)}-\Theta_{ij}^{(2)}\right)/\hat{\sigma}_{ij}^{(2)}$ among different methods.

### v3

@@ 3.2 Support recovery and asymptotic distribution
Assumption 4 (Minimum signal conditions for node $j$)
For a node $j\in[p]$, there exists a sufficiently large constant $C_{1}>0$ such that
$\min_{k\in[K]}\min_{i\in S_{j}^{(k)}}\left|\Theta^{(k)}_{i,j}\right|^{2}\geq$ $\frac{C_{1}}{n_{k}}\left(\frac{\log p}{s_{0,j}}+\log(s_{j}K)\right),$ (8)
$\min_{i\in S_{j}}\sum_{k\in[K]}\frac{n_{k}}{N}\left(\Theta^{(k)}_{i,j}\right)^{2}\geq$ $C_{1}\left(\frac{\log p+s_{0,j}\log(s_{j}K)}{N}\right).$ (9)
For a node $j\in[p]$, this assumption requires: (i) each nonzero per-graph edge $\Theta^{(k)}_{i,j}$ to exceed the per-graph noise level (8), and (ii) each neighbor $i\in S_{j}$ to have sufficiently large aggregated squared strength across graphs (9). They play complementary roles in guaranteeing exact edge- and neighbor-level support recovery. These two signal conditions are not implied by one another, because $s_{0,j}$ captures only the average occurrence for the neighbors across graphs, not the exact number of graphs in which a specific iji-j edge appears. Based on this assumption, Theorem 2 establishes the local selection consistency.
@@
[…]
@@
    Selection consistency: (^S_j^(1), , ^S_j^(K) ) = (S_j^(1), , S_j^(K) ).
@@
    Sharper estimation error bound: _k [K] nkN ∥^Θ^(k)_, j - Θ^(k)_, j _2^2 K + sjs0,j+ logpN.
Theorem 2 shows that our method is inherently signal-adaptive: it adaptively achieves support recovery and sharper estimation rates for those nodes with strong signals, and maintains the results in Theorem 1 for those nodes with relatively weak signals. This property is a distinct advantage of our dynamic regularized method over the existing joint estimation methods Ma and Michailidis (2016); Cai et al. (2016). Furthermore, define $\hat{S}_{j+}^{(k)}:=\hat{S}_{j}^{(k)}\cup\{j\}=\text{supp}\left(\hat{\Theta}_{\cdot,j}^{(k)}\right)$. The following theorem establishes the asymptotic property of nodes with strong signals.
@@
Assume that all conditions in Theorem 2 hold, and that $\sqrt{n}\succ\max_{j\in\mathcal{A}}\left(s_{j}\log p\right)$. Then, as $n\to\infty$, for every $(j,k)\in\mathcal{A}\times[K]$, we have
$\max_{i\in\hat{S}_{j+}^{(k)}}~\sup_{w\in\mathbb{R}}\left|\mathbf{P}\left(\frac{\sqrt{n_{k}}\left(\hat{\Theta}^{(k)}_{i,j}-\Theta^{(k)}_{i,j}\right)}{\sigma_{i,\hat{S}_{j}}^{(k)}}\leq w\right)-\Phi(w)\right|\to 0,$
@@
Theorem 3 demonstrates that our MIGHT method possesses asymptotic normality, thereby enabling statistical inference. This property further strengthens the signal adaptivity of the MIGHT estimator, demonstrating its superiority of our method. In application, we estimate $\left(\sigma^{(k)}_{i,\hat{S}_{j}}\right)^{2}$ by using
$\frac{1}{n_{k}}\sum_{\ell=1}^{n_{k}}\left\{\left(\hat{\Sigma}^{(k)}_{\hat{S}_{j+}^{(k)},\hat{S}_{j+}^{(k)}}\right)^{-1}_{i,\cdot}X_{\ell,\hat{S}_{j+}^{(k)}}^{(k)}\left(X_{\ell,\hat{S}_{j+}^{(k)}}^{(k)}\right)^{\top}\left(\hat{\Sigma}^{(k)}_{\hat{S}_{j+}^{(k)},\hat{S}_{j+}^{(k)}}\right)^{-1}_{\cdot,j}\right\}^{2}-\left(\hat{\Theta}_{i,j}^{(k)}\right)^{2},$
@@ 4 Numerical experiments
This section evaluates the performance of our proposed method, MIGHT, against five existing methods: three joint estimation methods: the joint estimator method by Guo et al. (2011) (JEM), the group graphical lasso by Danaher et al. (2014) (GGL), and the fast and scalable joint estimator by Wang et al. (2017) (FJEM); and two separate estimation methods: the separate graphical lasso by Friedman et al. (2008) (Sep Glasso) and the separate nodewise $L_{0}$-regression by Shu et al. (2024) (Sep Node). We use the Bayesian Information Criterion (BIC) to select the optimal hyperparameters for these five methods. All code for this section has been uploaded to https://github.com/ShixiangLIUstat/MIGHT-Algorithm.
We set $p=100,K=10$ and $n_{1}=\cdots=n_{K}=100$. Data generation begins with an Erdős–Rćenyi base graph $G$, where each pair of nodes connects independently with a probability of $0.1$. We generate $K$ graphs $G_{1},\cdots,G_{K}$ by independently pruning a fraction $\rho$ of edges from the base graph $G$, with $\rho\in(0,1)$ controlling neighbor-similarity and edge-sparsity. The value of each edge in $G_{k}$ is uniformly sampled from $1,-0.5]\cup[0.5,1]$, forming the adjacency matrix $\Omega^{(k)}$. Then, the positive-definite precision matrix $\Theta^{(k)}$ is constructed as $\Theta^{(k)}=\Omega^{(k)}+\left(r+|\Lambda_{\min}(\Omega^{(k)})|\right)\cdot\mathbf{I}_{p}$, where parameter $r>0$ modulates the signal strength by inflating diagonal entries, thereby reducing the signal-to-noise ratio (see (3)) in the nodewise regression. Finally, by applying the corresponding covariance matrices $\left\{\Sigma^{(k)}\right\}_{k\in[K]}=\left\{(\Theta^{(k)})^{-1}\right\}_{k\in[K]}$, we get $K$ datasets as $X^{(1)},\cdots,X^{(K)}$.
@@ 4.1 Simulation 1: similarity and sparsity
We fix $r=0.1$ and vary $\rho\in\{0.2,0.5,0.8\}$, with each setting 100 times simulation. Table 1 presents the simulation results; all MCC values have been multiplied by 100, and standard errors are shown in parentheses. Our MIGHT method demonstrates superior performance across all evaluation metrics compared to competing methods. As $\rho$ increases, the recurrence frequency of shared neighbors across the $K$ graphical models decreases, leading to diminished structural similarity and a decline in the MCC-Ngbr metric. Additional performance metrics (e.g., Hamming Loss and F1 score) are provided in Appendix F in the supplementary material.
@@
Estimation Accuracy Support Recovery
Methods Frobenius Norm Max L2 Norm MCC-Edge MCC-Ngbr
$\rho=0.2$
MIGHT 18.979 (0.113) 6.257 (0.088) 63.662 (0.305) 89.673 (0.267)
@@ 4.2 Simulation 2: signal strength
We fix $\rho=0.5$ and vary $r\in\{0.2,0.25,0.33,0.5,1\}$, with $1/r\in\{1,2,3,4,5\}$ measuring the signal strength. The results are presented in Figure 2. As the signal strength increases, the MIGHT method demonstrates lower estimation error and higher selection accuracy, confirming that enhanced signal strength contributes to better performance, consistent with our theoretical results.
Figure 2: Performance metrics with increasing signal strength ($1/r$), with each point $50$ times repetitions. All MCC values have been multiplied by 100. FJEM exhibits unstable estimation, and we exclude its excessively large estimation errors.
@@ 4.3 Simulation 3: asymptotic distribution
We fix $p=50$, $\rho=0.5$, $r=0.1$, and compare the asymptotic performance of MIGHT, GGL, JEM, and Sep Node methods across six specific matrix entries: three diagonal elements (Figure 3) and three off-diagonal elements (Figure 4). The results demonstrate that the MIGHT method achieves better asymptotic normality compared to the separate $\ell_{0}$ estimator (Sep Node), while GGL and JEM, two joint estimation methods based on M-estimator, fail to exhibit this property.
Figure 3: Histogram of (diagonal) $\sqrt{n}\left(\hat{\Theta}_{ii}^{(k)}-\Theta_{ii}^{(k)}\right)/\hat{\sigma}_{ii}^{(k)}$ among different methods with $300$ times repetitions. The orange curve represents the density of the standard normal distribution, while the black and orange vertical lines indicate the empirical mean and the population mean (0), respectively.
Figure 4: Histogram of (off-diagonal) $\sqrt{n}\left(\hat{\Theta}_{ij}^{(k)}-\Theta_{ij}^{(k)}\right)/\hat{\sigma}_{ij}^{(k)}$ among different methods with $300$ times repetitions.
@@ 5 Real data analysis
In this section, we apply our MIGHT method to a gene-expression cancer RNA-seq dataset, which is accessible via the UCI Machine Learning Repository Fiorini (2016). This dataset encompasses 801 patients and comprises 20,531 genes. The patients are categorized into five distinct types based on various tumor types, namely BReast CAncer (BRCA), KIdney Renal cell Carcinoma (KIRC), COlon ADenocarcinoma (COAD), LUng ADenocarcinoma (LUAD), and PRostate ADenocarcinoma (PRAD). The numbers of patients in these subtypes are 300, 78, 146, 141, and 136, respectively. Although these subtypes exhibit biological differences, they all involve tumor tissue, which necessitates the joint graphical model estimation.
@@ 5.2 The classification tasks based on QDA
$\delta_{k}(\mathbf{x})=\log\hat{\pi}_{k}+\frac{1}{2}\log\left|\hat{\Theta}^{(k)}\right|-\frac{1}{2}(\mathbf{x}-\hat{{\mu}}^{(k)})^{\top}\hat{\Theta}^{(k)}(\mathbf{x}-\hat{{\mu}}^{(k)}),$ (10)
where $k=1,\cdots,5$, $\hat{\pi}_{k}=n_{k}/N$ is the proportion of observations belonging to the class $k$, and ${\hat{\mu}}_{k}=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\bm{x}_{i}^{(k)}$ is the empirical mean for the training set of each class $k$. The classification rule is then given by $\hat{k}(\mathbf{x})=\arg\max_{k=1,\cdots,5}\delta_{k}(\mathbf{x})$.
@@ 6 Conclusion and discussion
This paper proposes an adaptive and computationally efficient method for the joint estimation of multiple graphical models. Theoretical analysis demonstrates that the proposed method achieves superior estimation accuracy and establishes selection consistency and asymptotic property. Both simulation studies and empirical analyses validate the superiority of our method.
@@
Moreover, our joint-learning framework can be well-suited to modern applications. For example, in the cross-domain recommendation task, a user’s interactions with the same items across different domains are often correlated. Joint estimation can effectively identify these similar structures across domains, enabling cross-domain knowledge fusion through multi-graph neural networks Ouyang et al. (2021). Moreover, our framework shows broader applicability in multiple directed acyclic graphs, multimodal data analysis, and so on. We leave them for future exploration.
[…]
Appendix A Additional simulation results
This section presents supplementary results for the simulation studies in Section 4.1. All parameters are the same as in Section 4.1. In terms of estimation accuracy, we additionally evaluate the maximum column-wise $L_{1}$-norm (Max L1 Norm), defined as $\max_{j\in[p]}\sum_{k\in[K]}\frac{n_{k}}{N}\left\|\hat{\Theta}^{(k)}_{\cdot j}-\Theta^{(k)}_{\cdot j}\right\|_{1}$. Regarding support set recovery, we incorporate Precision, Recall, F1 score, and Hamming Loss, calculated as follows:
$\text{Precision}=$ $\frac{TP}{TP+FP},$
$\text{Recall}=$ $\frac{TP}{TP+FN},$
$\text{F1 score}=$ $\frac{2}{(\text{Precision})^{-1}+(\text{Recall})^{-1}},$
$\text{Hamming loss}=$ $FP+FN,$
where TP denotes true positives, FP denotes false positives, and FN denotes false negatives. We compute these four quantities at two levels: (i) edge-level selection accuracy across all $K$ graphs (reported as Precision-Edge, Recall-Edge, F1-Edge, Hamming-Edge), and (ii) neighbor-level selection accuracy using the union $S_{j}=\bigcup_{k\in[K]}S_{j}^{(k)}$ for each node $j\in[p]$ (reported as Precision-Ngbr, Recall-Ngbr, F1-Ngbr, Hamming-Ngbr). All reported values are averages over 100 times Monte Carlo replications. Higher Precision/Recall/F1 and lower Max L1 Norm/Hamming loss indicate better performance.
Table 3 presents the supplementary results. All Precision/Recall/F1 values have been multiplied by 100, and standard errors are shown in parentheses. Consistent with Table I in Section IV-A, across all examined values of $\rho$, our MIGHT method consistently attains superior estimation accuracy compared to competing methods. In terms of the edge-level and the neighbor-level support recovery, the MIGHT estimator preserves relatively high Precision, and attains the best overall performance on composite metrics such as F1 score and Hamming loss.
Table 3: Additional measurements across varying $\rho$.
Estimation Edge-wise Support Recovery Neighbor-wise Support Recovery
Methods Max L1 Norm Precision-Edge Recall-Edge F1-Edge Hamming-Edge Precision-Ngbr Recall-Ngbr F1-Ngbr Hamming-Ngbr
$\rho=0.2$
MIGHT 8.023 (0.112) 98.266 (0.068) 44.479 (0.443) 61.202 (0.424) 4412.000 (68.402) 98.797 (0.096) 83.739 (0.547) 90.619 (0.327) 170.067 (6.506)
GGL 10.588 (0.124) 49.007 (0.542) 53.213 (0.496) 50.926 (0.290) 8029.333 (112.060) 25.086 (1.176) 97.219 (0.980) 39.294 (1.058) 3040.933 (109.350)
JEM 9.527 (0.147) 94.766 (0.146) 36.176 (0.656) 52.268 (0.703) 5155.933 (83.513) 99.855 (0.040) 52.077 (0.800) 68.347 (0.700) 470.800 (10.502)
FJEM 310.986 (12.031) 8.130 (0.069) 95.529 (0.241) 14.983 (0.117) 84806.800 (299.661) 9.888 (0.086) 100.000 (0.000) 17.993 (0.143) 8921.067 (8.536)
Sep Glasso 13.169 (0.130) 48.674 (0.304) 24.573 (0.335) 32.618 (0.314) 7932.200 (75.785) 33.437 (0.311) 78.050 (0.533) 46.768 (0.283) 1740.800 (23.391)
Sep Node 11.835 (0.138) 96.191 (0.262) 5.954 (0.089) 11.210 (0.159) 7377.600 (69.475) 95.383 (0.305) 37.898 (0.459) 54.203 (0.492) 626.533 (9.288)
$\rho=0.5$
MIGHT 4.645 (0.089) 96.415 (0.088) 54.565 (0.634) 69.628 (0.521) 2341.600 (44.850) 98.776 (0.099) 79.159 (0.588) 87.850 (0.357) 215.800 (6.790)
GGL 6.917 (0.123) 52.231 (0.653) 67.697 (0.917) 58.802 (0.523) 4675.267 (95.499) 45.193 (1.611) 87.357 (1.218) 58.645 (1.412) 1301.933 (112.747)
JEM 6.142 (0.096) 88.103 (0.230) 42.651 (0.991) 57.292 (0.903) 3112.267 (60.960) 99.752 (0.074) 53.247 (0.943) 69.293 (0.805) 462.733 (11.435)
FJEM 305.408 (16.413) 5.180 (0.042) 96.528 (0.231) 9.831 (0.075) 87189.600 (392.454) 9.949 (0.078) 100.000 (0.000) 18.095 (0.128) 8915.067 (7.679)
Sep Glasso 8.716 (0.119) 49.300 (0.323) 37.201 (0.622) 42.309 (0.404) 4982.133 (51.580) 35.931 (0.345) 83.046 (0.634) 50.094 (0.311) 1633.600 (27.844)
Sep Node 7.340 (0.139) 97.050 (0.225) 10.790 (0.217) 19.401 (0.351) 4409.733 (41.578) 96.462 (0.278) 42.081 (0.688) 58.508 (0.672) 586.533 (10.120)
$\rho=0.8$
MIGHT 1.767 (0.032) 94.462 (0.184) 70.161 (0.496) 80.488 (0.335) 673.800 (12.248) 97.949 (0.148) 73.731 (0.460) 84.106 (0.294) 245.933 (4.865)
GGL 2.786 (0.041) 37.528 (0.486) 87.525 (0.429) 52.441 (0.432) 3163.067 (68.096) 45.296 (0.398) 87.855 (0.493) 59.707 (0.298) 1048.133 (14.790)
JEM 2.582 (0.037) 83.763 (0.404) 60.347 (0.774) 70.029 (0.449) 1020.933 (14.343) 99.659 (0.069) 59.727 (0.772) 74.607 (0.594) 357.533 (7.459)
FJEM 513.932 (36.266) 2.084 (0.019) 98.942 (0.172) 4.082 (0.037) 92248.667 (469.228) 8.918 (0.062) 100.000 (0.000) 16.374 (0.105) 9017.133 (6.150)
Sep Glasso 3.708 (0.044) 46.726 (0.306) 68.719 (0.602) 55.566 (0.243) 2178.333 (25.182) 37.463 (0.284) 87.771 (0.464) 52.471 (0.243) 1405.933 (19.041)
Sep Node 2.581 (0.039) 94.732 (0.312) 31.719 (0.382) 47.491 (0.442) 1388.600 (13.522) 93.738 (0.380) 53.858 (0.575) 68.365 (0.511) 439.267 (6.891)